英語 (English)

主題別英語 (Thematic English) (総科 ((人文・社創・総理 B)1 年))

フロストダン · 非常勤講師 1 単位後期火 3 ・ 4

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 To present topical themes encapsulating grammatical patterns. This encourage the student to see modern global values and see English as an important tool in understanding them. It will also aid their ability to respond appropriately to such ideas.

【授業の概要】 The course will consist of a series of themes listed below, which will be explored. Students will be required to produce works on those themes to be presented verbally and in writing to enforce the stated goal. These exercises will be derived from handouts as well as from multimedia presentations. Progression through the themes will depend on the student’s ability but hopefully one unit will take two classes.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】

1. Through the use of ESL activities build up the skill and confidence level of the students

2. Use group work to promote better exploration of the themes

3. Require presentation of work completed to ensure real understanding of the

subject matter covered. 【授業の計画】

1. Nations and Nationalities around the world

2. Brands and Advertising

3. Great People

4. Society, traditions, and problems

5. Dreams, Hopes, and Ambitions 【教科書】 Handouts and other class materials will be given 【参考書等】 [参考資料] 【成績評価の方法】 Evaluation is based on attendance, participation in class, completion

of homework, as well as examinations.

【再試験の有無】 Consideration will be given for reexamination. Meritorious work will also be accepted in lieu of missed assignments. Individual performance will also affect the final grade

【受講者へのメッセージ】 Please always bring dictionary, notebook, pen, pencil, and red pencil

to class. Pocket translators can be used.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221149 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ フロスト .

English

Thematic English (総科 ((人文・社創・総理 B)1 年)) Dan Frost · Part-time Lecturer 1 unit 後期火 3 ・ 4

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ To present topical themes encapsulating grammatical patterns. This encourage the student to see modern global values and see English as an important tool in understanding them. It will also aid their ability to respond appropriately to such ideas.

Outline⟩ The course will consist of a series of themes listed below, which will be explored. Students will be required to produce works on those themes to be presented verbally and in writing to enforce the stated goal. These exercises will be derived from handouts as well as from multimedia presentations. Progression through the themes will depend on the student’s ability but hopefully one unit will take two classes.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩

1. Through the use of ESL activities build up the skill and confidence level of the students

2. Use group work to promote better exploration of the themes

3. Require presentation of work completed to ensure real understanding of the

subject matter covered.

Schedule⟩

1. Nations and Nationalities around the world

2. Brands and Advertising

3. Great People

4. Society, traditions, and problems

5. Dreams, Hopes, and Ambitions Textbook⟩ Handouts and other class materials will be given Reference⟩ [参考資料] Evaluation Criteria⟩ Evaluation is based on attendance, participation in class,

completion of homework, as well as examinations.

Re-evaluation⟩ Consideration will be given for reexamination. Meritorious work will also be accepted in lieu of missed assignments. Individual performance will also affect the final grade

Message⟩ Please always bring dictionary, notebook, pen, pencil, and red pencil

to class. Pocket translators can be used.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221149 Contact (Office-Hour, Room, E-mail)⟩

⇒ Frost .